# OATS: Outlier-Aware Pruning through Sparse and Low Rank Decomposition

### v2

@@ 3.4 Hardware Speedup
We benchmark, using the DeepSparse engine by NeuralMagic (2021), the CPU throughput induced by OATS compared to models pruned with unstructured sparsity. We run end-to-end inference on a compressed Phi-3 Medium 15B model for a single batch of 2048 tokens on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores. The achieved throughput and speedup (over a dense model) are shown in Table 7 below. By trading unstructured sparsity for structured sparsity through the low-rank terms, OATS achieves greater CPU speed-up compared to methods that rely solely on unstructured pruning. Notably, at $40\%$ compression, OATS is 1.37×1.37$\times$1.37 × faster than unstructured pruning.
@@ A.5 Magnitude-Based Pruning for the Sparse Component
A.6 Additional Hyperparameter Tests for OATS
@@
A.7 Performance Gap Between OATS and Wanda
@@ A.9 MMLU Subjects
A.10 Attention Rollout: Details
@@
A.11 Zero-Shot Task-Specific Performance
@@ A.12 Wikitext-2 Perplexity
A.13 Pruning Implementation and Hyperparameters

### v3

@@ 3.4 Hardware Speedup
We benchmark, using the DeepSparse engine by NeuralMagic (2021), the CPU throughput induced by OATS compared to models pruned with unstructured sparsity. We run end-to-end inference on a compressed Phi-3 Medium 15B model for a single token on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores (for higher token-counts see Appendix A.6). The achieved throughput and speedup (over a dense model) are shown in Table 7 below. By trading unstructured sparsity for structured sparsity through the low-rank terms, OATS achieves greater CPU speed-up compared to methods that rely solely on unstructured pruning. Notably, at $40\%$ compression, OATS is 1.37×1.37$\times$1.37 × faster than unstructured pruning.
@@ A.5 Magnitude-Based Pruning for the Sparse Component
A.6 Additional CPU Speedup Experiments
We provide additional experiments measuring the CPU speedup of OATS compared to dense models. Following the same setup as Section 3.4, we run end-to-end inference on a compressed Phi-3 Medium 15B model, but instead with a sequence consisting of 256 tokens, on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores. The results are presented in Table 14 below:
Compression Method Throughput Speedup
0% Dense 0.37 1.00×1.$0$
30% Unstructured Pruning 0.40 1.08×1.08$\times$1.08 ×
OATS 0.40 1.08×1.08$\times$1.08 ×
40% Unstructured Pruning 0.43 1.16×1.16$\times$1.16 ×
OATS 0.43 1.16×1.16$\times$1.16 ×
50% Unstructured Pruning 0.49 1.32×1.32$\times$1.32 ×
OATS 0.49 1.32×1.32$\times$1.32 ×
Table 14: Comparison of throughput (tokens/second) and speedup achieved through OATS and unstructured pruning methods relative to their dense counterparts on a sequence consisting of 256 tokens.
Although OATS and unstructured pruning now deliver comparable CPU speed‑ups, we expect further gains are achievable by designing optimizations that explicitly exploit OATS’ sparse plus low‑rank structure.
A.7 Additional Hyperparameter Tests for OATS
@@
A.8 Performance Gap Between OATS and Wanda
@@ A.10 MMLU Subjects
A.11 Attention Rollout: Details
@@
A.12 Zero-Shot Task-Specific Performance
@@ A.13 Wikitext-2 Perplexity
A.14 Pruning Implementation and Hyperparameters